OpenStudy (anonymous):

What is the equation of the line in standard form that passes through the point (–2, –3) and is parallel to the line y = x + 9?

OpenStudy (anonymous):

Well, we know the equation for a line is \(y=mx+b\) Where \(m\) is its slope and \(b\) is the y-intercept. We also know that if Line A has a slope of \(m_1\) and Line B has a slope of \(m_2\), and Line A is parallel to Line B, then \(m_1=m_2\). So we can develop some things from here. One thing is that we know the mentioned line, has a formula of \(y=x+9\); from that, we can deduce that it has a slope of \(1\) and therefore, our line must also have a slope of \(1\) since they are parallel

OpenStudy (anonymous):

So the formula for our line is: \[y=(1)x+b=x+b\] We can sub in a point we know: \[-3=-2+b\] And solve for \(b\) \[b=-3+2=-1\] So the formula would be \[y=x-1\]

OpenStudy (anonymous):

A. x – y = 1 B. x + y = 1 C. x – y = –5 D. x + y = –5 are my answers to choose from

OpenStudy (anonymous):

First, find the slope of the equation of \[y = x + 9\] by using your slope-intercep form: \[y = mx + b\]. In this case, m = 1. Using the point slope formula \[(y-y1) = m(x-x1)\] Use this formula to plug in the numbers that have been provided. Points (-2,-3) and m = 1. Then solve for y.

OpenStudy (anonymous):

Oh okay well, \[y=x-1\] \[y-x=-1\] \[-(y-x)=1\] \[x-y=1\] So the equation is A

OpenStudy (anonymous):

Thank you bunches!! =)

OpenStudy (anonymous):

Anytime :)